I am trying to figure out how old my great grandma,my grandma,and my mom are.mom gave me some clues.She said ,I am four times older than you ,Samantha,Grandma,my mom is 24 years older than me.My mother,your grandma is forty years younger than her mom,your great grandma.How old is Samantha,,Mom,Grandma, and great grandma
Found 3 solutions by mananth, josgarithmetic, ikleyn:
Answer by mananth(16946)


She said ,I am four times older than you ,Samantha,Grandma,my mom is 24 years older than me.My mother,your grandma is forty years younger than her mom,your great grandma.How old is Samantha,,Mom,Grandma, and great grandma
let my age be x years
mom's age =4x
Grandma = 4x+24
Great grand ma = 4x+24+40
any other data to find ages?

Answer by josgarithmetic(39618)

Description is not enough.

age p for the person making the description
age v for the great-grandmother
PEOPLE              AGE
First Person         p
Mother               p+4p=5p
Grandmother         24+(5p)  and  24+5p=v-40
Greatgrandmother     v

Not possible to exactly find age of any of them.

Notice there is ONE equation in two variables. A reasonable example for p could be p=5 for First Person. This will give:
Mother is 25
Grand Mother is 49
GreatGrand Mother is 79

Answer by ikleyn(52797)

.

It NEVER was so many nonsensical posts as today.

Where do you recover these nonsensical problems, that have no solutions and CAN NOT be solved ?